OpenStudy (anonymous):

how can i integrate cos2t without using substituion

OpenStudy (anonymous):

Realize that if you differntiate it you have to do the chain rule giving -2sin(2t) So that means if you integrate it you need a factor of (1/2) to take care of the 2 outside. So the integral of cos(2t)=(1/2)sin(2t) Because: \[\frac{d}{dt}(\frac{1}{2})\sin(2t)=\cos(2t)\]

OpenStudy (anonymous):

why is it -2sin(2t), why a negative

OpenStudy (anonymous):

Because the derivative of cos(x)=-sin(x)

OpenStudy (anonymous):

The negative isn't too important You can use hyperbolic trig functions (they don't have the negative) \[\frac{d}{dt}\sinh(3t)=3\cosh(3t) \rightarrow \int\limits \frac{d}{dt}\sinh(3t)=3\int\limits \cosh(3t) \rightarrow \frac{1}{3}\sinh(3t)=\int\limits \cosh(3t)\] You just need the fraction to compensate for the 3t inside the function ("undoing" the chain rule, if you will).